### v1

# UrbanVLP: A Multi-Granularity Vision-Language Pre-Trained Foundation Model for Urban Indicator Prediction

Abstract.
Urban indicator prediction aims to infer socio-economic metrics in diverse urban landscapes using data-driven methods. However, prevalent pre-trained models, particularly those reliant on satellite imagery, face dual challenges. Firstly, concentrating solely on macro-level patterns from satellite data may introduce bias, lacking nuanced details at micro levels, such as architectural details at a place. Secondly, the lack of interpretability in pre-trained models limits their utility in providing transparent evidence for urban planning. In response to these issues, we devise a novel Vision-Language Pre-Trained Model (UrbanVLP) in this paper. Our UrbanVLP seamlessly integrates multi-granularity information from both macro (satellite) and micro (street-view) levels, overcoming the limitations of prior pre-trained models. Moreover, it introduces automatic text generation and calibration, elevating interpretability in downstream applications by producing high-quality text descriptions of urban imagery. Rigorous experiments conducted across six socio-economic tasks underscore UrbanVLP’s superior performance. We also deploy a web platform to verify its practicality.
Y. Liang is the corresponding author. Email: yuxliang@outlook.com
1. Introduction
Urban Indicator Prediction, a scholarly pursuit in the realm of urban computing (Zheng et al., 2014), deploys data-driven methodologies to forecast socio-economic metrics (e.g., GDP, population, carbon emission) in diverse urban landscapes. Rooted in the escalating global prominence of urban environments and the imperative for judicious urban planning, this academic discipline finds motivation in enhancing the efficacy of policymaking, optimizing resource allocation, and mitigating challenges endemic to urban agglomerations (Xi et al., 2022; Yan et al., 2023).
Classical machine learning approaches commonly tackle urban indicator prediction via supervised training, leveraging task-specific datasets like satellite imagery. These tasks span a spectrum from predicting poverty levels (Jean et al., 2016a; Han et al., 2020; Perez et al., 2017; Ayush et al., 2021a), crop yields (You et al., 2017; Wang et al., 2018; Martinez et al., 2021; Yeh et al., 2021), land cover (Uzkent et al., 2019), to commercial activity (He et al., 2018; Liu et al., 2023d). Nevertheless, the inherent task specificity of supervised learning, reliant on ample labeled data, may impede the model’s capacity for broad generalization. This constraint poses a conceivable risk to the overarching robustness and effectiveness of the model.
In contrast, Pre-trained Model, also known as Self-Supervised Learning (Jaiswal et al., 2020), empowers the autonomous discovery and optimization of urban region representations from diverse urban data modalities, encompassing satellite images (Jean et al., 2016b; Ayush et al., 2021b; Yao et al., 2022b; Watmough et al., 2019), Points of Interests (POIs) (Fu et al., 2019; Jenkins et al., 2019; Zhang et al., 2021; Li et al., [n. d.]), road networks (Liang et al., 2018, 2021; Wei Chen, 2023) and human mobility (Wang and Li, 2017; Seong et al., 2023; Li et al., 2023b, a). By obviating the necessity for externally labeled datasets, these pre-training techniques enable models to learn intrinsic region representations, fostering adaptability and scalability in comprehending intricate urban dynamics for enhanced generalization across various urban indicator prediction tasks.
Figure 1. Single-granularity vs. Multi-granularity modeling.
Given the widespread accessibility of satellite imagery through platforms such as Google Maps or Baidu Maps (Li et al., 2022b; Liu et al., 2023d; Xi et al., 2022), coupled with its wealth of information on regional features (e.g., road network patterns, building density, vegetation coverage), a predominant majority of initiatives harness satellite imagery [^1] As other modalities (e.g., mobility, POIs) are always inaccessible due to privacy issues. as the foundational modality for learning urban representations. Despite its promise, urban environments exhibit a spatial hierarchy in reality, from a macro region level to a micro location level (e.g., architectural details, street furniture, storefront). As depicted in Figure 1, while the satellite images (a) and (b) possess remarkably similar layouts, their urban indicators considerably differ. Upon zooming into micro levels using the corresponding street-view images (c) and (d), a more nuanced and fine-grained understanding becomes apparent. Exclusively capturing macro-level patterns through satellite imagery may introduce substantial bias, falling short in furnishing more comprehensive information across diverse urban levels. Hence, the integration and alignment of multi-granularity information (macro-level and micro-level) for urban indicator prediction present an ongoing challenge that requires further exploration.
Furthermore, while urban decision-makers rely on explainable evidence for service deployment and urban planning, the pre-trained models for urban indicator prediction mostly lack interpretability. They typically learn direct mapping from single or multi-modal data to the output representations, which poses a significant challenge in comprehending the intrinsic factors influencing the predictive outcomes. In light of the recent surge in Large Language Models (LLMs) (Ouyang et al., 2022; Zhao et al., 2023), UrbanCLIP (Yan et al., 2023) emerges as an innovative solution. This model effectively integrates textual modality as auxiliary knowledge into the realm of urban indicator prediction. Figure 1 shows how these textual data encapsulate key elements within both satellite images and their street-view counterparts. Notably, even in the absence of image-text pairs, UrbanCLIP adeptly harnesses a myriad of textual descriptions generated by pre-trained image-to-text LLMs, thereby facilitating both explainable reasoning and prediction. However, the text generation process of UrbanCLIP raises several critical concerns. a) Hallucination: The generation of textual content sometimes deviates from or introduces information not present in the input satellite imagery. b) Homogenization: The generated description appears overly simplified and general, potentially leading to homogenization and inadequate differentiation. For instance, the average length of the text description in UrbanCLIP is quantified at 18.1 words, leading to insufficient information. To better align with the intricacies of satellite (or street-view) imagery, we necessitate a more powerful approach that goes beyond the intuitive text generation method in UrbanCLIP, ensuring a more faithful and explainable representation of urban regions.
In this paper, we present a Vision-Language Pre-Trained Model entitled UrbanVLP for urban indicator prediction. Our model elaborately integrates multi-granularity (macro and micro-level) information from both satellite and street-view imagery to produce comprehensive urban region representations, while simultaneously harnessing the interpretability inherent in high-quality textual descriptions. Targeting the first challenge, we introduce a novel Multi-Granularity Cross-Modal Alignment module. It utilizes dual-branch contrastive learning to establish alignment between information derived from two semantic granularities. Compared to standard contrastive learning applied in the satellite branch at the macro level, we employ token-level contrastive learning to capture fine-grained details in street-view imagery. Concurrently, spatial information, enriched with the location coordinates of street-view images, is encoded and injected to provide essential spatial context. Finally, the fusion process aggregates features from street-view images and location-specific features within the same region, seamlessly incorporating fine-grained information into the satellite branch.
To address the second issue (i.e., lack of interpretability), we devise an Automatic Text Generation together with a Calibration mechanism to uphold text quality standards. Given the impracticality of generating a large number of descriptions using GPT-4V (Yang et al., 2023) or Gemini (Fu et al., 2023) due to prohibitive expense, we utilize a closely comparable GPT4V (Yang et al., 2023) model called ShareGPT4V (Chen et al., 2023a), which is a strong caption model trained on 100k GPT4V (Yang et al., 2023) generated captions. To guarantee the quality of LLM-generated descriptions, we propose a reference-free metric PerceptionScore, motivated by the human evaluation system. This metric comprises two components: Clipscore and Cyclescore, working collaborateively to capture both semantic similarity and detail inclusiveness.
Overall, the main contributions of our work are summarized as:
  • Multi-granularity cross-modal alignment. We explored the role of two different visual data modalities at different semantic granularities - street-view and satellite imagery. We introduced street-view data to inject fine-grained semantic information based on the satellite imagery and provided a framework.
  • Automatic text generation and calibration. Powered by image-to-text LLMs, we generate text descriptions and implement a robust evaluation mechanism, ensuring the fidelity between the generated texts and the corresponding image content.
  • A new benchmark & Empirical evidence. We plan to open-source the first vision-text and multi-granularity urban dataset upon paper notification, consisting of six downstream tasks across the socio-economy. Extensive experiments are conducted on them. As illustrated in Figure 2, UrbanVLP outperforms existing approaches by an average improvement of 3.55% on the $R^{2}$ metric.
Figure 2. $R^{2}$ performance on Beijing and Shenzhen dataset.
Figure 3. Overall framework of our proposed UrbanVLP.
2. Preliminaries
2.1. Formulation
Definition 1 (Urban Region) We follow previous research (Xi et al., 2022; Yan et al., 2023) and evenly divide the target area (e.g. a city) into $L$ urban regions.
Definition 2 (Urban Imagery) Diverse visual images, including satellite and street-view images, facilitate an intuitive analysis of urban regions. Specifically, satellite images provide a coarse-grained overview of the structural composition of a region. Each input satellite image w.r.t the urban area $g$ can be represented as $I^{sa}_{g}\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ denote the length and width. Street-view images are captured by vehicles from map providers during urban traversals, offering detailed visual information on the fine-grained aspects within a region. By querying the latitude and longitude range of satellite images, multiple street-view images are associated together w.r.t the urban region $g$ based on their locations $\mathcal{L}_{g}$. Each individual street-view image is represented as $I^{{sv}}_{g}\in\mathbb{R}^{H\times W\times 3}$.
Definition 3 (Text Description) Textual descriptions of urban area $g$ (including satellite text $T^{sa}_{g}$ and street-view text $T^{sv}_{g}$) provide precise means for urban area analysis. Such text can be manually generated or produced using image annotation tools. Particularly, leveraging advanced LLMs enables detailed and insightful analytical descriptions of given area images. Notably, incorporating spatial context (e.g. vegetation coverage and POI) significantly reflects the functional characteristics of the region.
Definition 4 (Urban Indicators) Urban indicators assess various aspects of a region, including population, economy, society, and public services, from multiple perspectives. A set of $L$ urban regions with $K$ indicators is represented as $\mathbf{Y}\in\mathbb{R}^{L\times K}$. In this paper, we comprehensively use population (count), GDP (in million Chinese Yuan), carbon emissions (in tons), night lights (intensity), house prices (in Chinese Yuan), and points of interest, i.ePOI, (count for different categories) as ground-truth urban indicators.
Problem Statement (Urban Region Representation) We next formally define the problem of urban region representation learning. Given a satellite images $I^{sa}_{g}$, a set of street-view images $\mathcal{I}^{sv}_{g}$ belonging to the target coverage, and their corresponding latitude and longitude pair $\mathcal{L}_{g}$. The objective is to employ a learning function $\mathcal{F}$ to map them to representation vectors $\mathbf{e}_{g}=\mathcal{F}(I^{sa}_{g},\mathcal{I}^{sv}_{g},\mathcal{L}_{g})$. The representation can then be further utilized to infer urban indicators $\mathbf{Y}_{g}$ for any given region $g$. We assume that during the training phase, automated additional text descriptions are available.
2.2. Related Work
2.2.1. Urban Region Representation Learning
In literature, many studies have learned task-specific region representations from various urban data, especially urban imagery due to its consistent updates and easy accessibility (Li et al., 2022c; Liu et al., 2023d; Xi et al., 2022). For example, Jean et al. (2016b) used nighttime light images as weak labels to learn area representation and assess regional economic development, while Urban2Vec (Wang et al., 2020) integrated street-view imagery and POI data to learn neighborhood embeddings. Some contrastive learning approaches like PG-SimCLR (Xi et al., 2022) and UrbanCLIP (Yan et al., 2023) have shown success in representing urban regions through satellite image. However, the aforementioned works exclusively considered a single type of urban imagery and overlooked the potential synergy between satellite imagery and street-view imagery, which can complement each other. Recently, some efforts (Barbierato et al., 2020; Huang et al., 2023a; Li et al., 2022c; Cao et al., 2023, 2018a; Law et al., 2019; Cao et al., 2018b; Huang et al., 2023b) have begun to explore the simple combination of satellite imagery and street-view imagery. For instance, Vision-LSTM (Huang et al., 2023b) integrates diverse street-level images into the satellite image domain using an LSTM-based network for precise urban village identification. Li et al. (2022c) propose utilizing street segments for region modeling and predict socio-economic indicators by merging street-view and satellite images through a multi-stage pipeline. Nevertheless, prior studies overlook distinct features in satellite and street-view images for varied granularities. Additionally, few studies incorporate text as a supervisory signal.
2.2.2. Large Multimodal Model
As LLMs undergo rapid evolution, a faction within the research community is increasingly focused on integrating visual knowledge into these models, such as CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021). Early studies (Li et al., 2022a, 2023c) enhanced CLIP with refined data strategies for more diverse datasets, proving effective for fundamental visual tasks (Li et al., 2022d; Liu et al., 2023c; Zhang et al., 2022) but less so for complex tasks like visual question answering. Recently, new models like MiniGPT-4 (Chen et al., 2023b), LLaVA (Liu et al., 2023b), and InstructBLIP (Dai et al., 2023) have improved the understanding of complex problems by expanding parameters and training data, called large multimodal models (LMMs). In this paper, we comprehensively evaluate the text generation quality of representative LMMs model and propose an automated text description generation and calibration technique to ensure effective image description capabilities with LMMs.
2.2.3. Vision-Language Pre-training (VLP)
VLP aims to jointly encode vision and language in a fusion model. Early works on VLP can be broadly categorized into single-stream (Chen et al., 2020; Kim et al., 2021; Li et al., 2020) and two-stream (Tan and Bansal, 2019; Lu et al., 2019) methodologies, with the former using a unified architecture for modality integration and the latter employing separate encoders before merging modalities. Milestone work CLIP (Radford et al., 2021) and its variants (Li et al., 2021; Yao et al., 2022a) highlight the efficacy of contrastive learning in cross-modal downstream tasks, such as zero-shot learning and cross-modal retrieval. Recent works (Tsimpoukelli et al., 2021; Alayrac et al., 2022) shift towards leveraging LLMs knowledge for vision-language representation learning. In urban indicator prediction, the potential of VLP paradigm remains untapped, with limited exploration of the benefits of textual information. Our groundbreaking work introduces multiple types of text descriptions as supplements, aiming to learn more comprehensive and interpretable urban region representations.
3. Methodology
We present an overview of our proposed framework in Figure 3. There are two principal stages in our framework:
  • Stage 1: We first devise an automatic text generation and calibration module. Leveraging ShareGPT4V (Chen et al., 2023a) (current state-of-the-art LMM captioner model), textual descriptions are created for each street-view image, incorporating geographical and visual prompts. PerceptionScore is also introduced to ensure high-quality text retention from both semantic and visual perspectives. Building upon the foundation of high-quality image-text pairs, we further develop a multi-granularity cross-modal alignment framework, achieving multi-granular alignment and multimodal correlation through multi-level contrastive learning and fine-grained information injection.
  • Stage 2: In the urban indicator prediction stage, we leverage the frozen encoder pre-trained in Stage 1 to extract input data features. Subsequently, a lightweight multi-layer perceptron (MLP) is fine-tuned to yield precise urban indicator predictions. Additionally, optional information such as street-view imagery and geographical coordinates is considered.
3.1. Automatic Text Generation and Calibration
Text Generation. For each satellite image, we initiate a range query to obtain the corresponding set of available street-view images and their respective geospatial coordinates within the designated area. Existing works (Lai et al., 2023; Fan et al., 2023a; Yan et al., 2023) proved that the quality of the text is crucial for enhancing performance. Therefore, advanced LMMs, owing to their robust cross-modal knowledge storage capabilities, offer detailed textual comprehension and descriptions for each street-view image. However, this confronts two primary challenges. Firstly, due to substantial financial implications associated with API usage, employing closed source LMMs like GPT4V (Yang et al., 2023) or Gemini (Fu et al., 2023) for generating tens of thousands of text descriptions is impractical. Secondly, street-view images often encompass diverse detailed information, necessitating a well-designed prompt template for obtaining high-quality textual descriptions.
To address these issues, we first utilize a closely related GPT4V model, known as ShareGPT4V (Chen et al., 2023a), a powerful language model with vision capabilities trained on 100k GPT4V generated captions, to generate street-view image descriptions. In the process of designing text prompt templates, the proportion of segmented elements in street-view images reflects the distribution of elements, aiding LMMs in evaluating the importance of each element. To achieve this, we employ a pre-trained segmentation model to decouple visual elements and calculate the proportion of segmentation for each element. Simultaneously, urban and geospatial coordinates of street-view images are included as prompts to guide the expression of additional urban information. Finally, the prompt template is formulated as follows: Analyze the street-view panoramic image in [City] in a comprehensive and detailed manner. The coordinate of the street-view image is [Longitude, Latitude]. The segmentation ratio of the street-view image is [Segment Information] . More prompt examples and segmentation ratio details can be found in Appendix C.
Text Calibration. In contrast to previous research (Yan et al., 2023) that employed simplistic rules and manual refinement processes for text description of satellite images, street view images often contain rich fine-grained information, leading to potential hallucination issues (Rawte et al., 2023) in generated text descriptions that require further calibration. Rule-based methods are overly general and fail to address specific issues in each image description as they are tailored to particular cases. On the other hand, manual rewriting is labor-intensive and demands significant effort. Therefore, an effective method for evaluating text quality is crucial.
[…]
$\begin{split}&I^{\prime}=\operatorname{SDXL}(I),\\ &\texttt{CycleScore}(I,I^{\prime})=1-\operatorname{MAE}(\operatorname{Seg}(I),\operatorname{Seg}(I^{\prime})),\\ &\texttt{PerceptionScore}(I,T)=(\texttt{CLIPScore}(I,T)+\mathop{\texttt{CycleScore}(I,I^{\prime})})/2.\end{split}$
Figure 4. The procedure of Cycle Score calculation.
3.2. Multi-Granularity Cross-modal Alignment
Modality Representation. Utilizing automated generation techniques, we obtain a dataset of high-quality image-text pairs, represented as $D=(I,T)$. Here, $I$ denotes either a satellite image $I^{sa}_{g}$ or a street view image $I^{sv}_{g}$, while $T$ represents either a satellite text description $T^{sa}_{g}$ or a street view text description $T^{st}_{g}$.
we first deploy Vision Transformer (ViT) (Dosovitskiy et al., 2021) as an image encoder to process images derived from satellites and street-views, proficient in deconstructing images into patches for feature extraction. In particular, we reshape the input image $I\in\mathbb{R}^{H\times W\times 3}$ into a sequence of flattened 2D patches ${I}_{P}\in\mathbb{R}^{N\times({P^{2}}\cdot{C})}$, where C is channel number, $(P,P)$ is the resolution of each image patch, and $N=HW/P^{2}$ is the patch number, which is the input sequence length of ViT. Notebly, the ViT utilizes a consistent latent vector size (denoted as $d$) across all its layers. Thus, we can obtain the outputs as the patch embeddings through a trainable linear projection. We also prepend a learnable embedding to the patch embedding sequence (i.e., $\mathbf{z}_{0}^{0}={I}_{cls}$). Besides, we add learnable position embedding $\mathbf{E}_{pos}$ to retain positional information. The process is as follows:
(1) $\mathbf{z}_{0}=[{I}_{cls};{I}_{p}^{1}\mathbf{W};{I}_{p}^{2}\mathbf{W};\cdots;{I}_{p}^{N}\mathbf{W}]+\mathbf{E}_{pos},$
where $\mathbf{W}\in\mathbb{R}^{(P^{2}\cdot C)\times d}$ and $\mathbf{E}_{pos}\in\mathbb{R}^{(N+1)\times d}$ are learnable parameter. $\mathbf{z}_{0}$ is the linear projection result and then sent to ViT.
Consistent with the standard Transformer encoder, ViT includes alternating layers of multi-head self-attention $\operatorname{MSA(\cdot)}$ operation (Vaswani et al., 2017). Furthermore, layer normalization (LN) is adopted before every block, whereas residual connections are after every block:
(2) $\mathbf{z}_{l}=\mathrm{MSA}(\mathrm{LN}(\mathbf{z}_{l-1}))+\mathbf{z}_{l-1},\quad l=1\ldots L.$
Finally, we use the output $\mathbf{z}_{l}^{0}$ corresponding to the special tokens ${I}_{cls}$ inserted at the beginning as the final output of the image encoder, rewrite as $\mathbf{z}_{I}$ for image representation.
Concurrently, we use basic Transformer-Encoder (Vaswani et al., 2017) as a text encoder by default to comprehend and encode the semantics of textual descriptions in parallel. The text encoder follows a similar $\operatorname{MSA}$ mechanism of the visual encoder. The input text sequence is bracketed with $[SOS]$ and $[EOS]$ tokens, and the activation of the highest layer of Transformer at $[EOS]$ token is considered the global representation $\mathbf{z}_{T}$ of text.
Moreover, depending on whether it is a satellite image-text pair or a street-view image-text pair, we categorize them into two branches with separate visual and text encoders – namely, the satellite branch and the street-view branch.
Furthermore, considering the longitude and latitude information corresponding to street view images, which aids in identifying the geospatial location of the area, we leverage the open-source GeoCLIP’s (Vivanco et al., 2023) location encoder. This encoder captures longitude and latitude information at different resolutions, generating semantically rich high-dimensional features. Here, a set of longitude and latitude $\mathcal{L}_{g}$ is aggregated into a single vector, denoted as $\mathbf{z}_{L}$, thereby incorporating geographical positioning.
Modality Alignment. Our goal is to capture the coarse-grained global regional information from the satellite branch and the fine-grained local positional information from the street-view branch.
For capturing global information, even though we have obtained overall visual features for the region from images in the satellite branch, we can further incorporate fine-grained information from images in the street-view branch. Specifically, the visual features of streets within each satellite region are aggregated, simultaneously fused with the corresponding geographical feature, aiding in learning a robust and comprehensive representation of urban areas. The formalized process is as follows:
(3) $\mathbf{z}_{g_{sa}}=\textit{f}~{}(\mathbf{z}_{I_{sa}},~{}\operatorname{Aggr}(\mathbf{z}_{I_{sv}}^{1},\mathbf{z}_{I_{sv}}^{2},...,\mathbf{z}_{I_{sv}}^{m}),~{}\operatorname{Aggr}(\mathbf{z}_{L}^{1},\mathbf{z}_{L}^{2},...,\mathbf{z}_{L}^{m})),$
where $\mathbf{z}_{I_{sa}}$ represents the satellite visual features for the queried region, $\mathbf{z}_{I_{sv}}^{k}$ denotes the $k$-th street-view visual feature in that region, and $\mathbf{z}_{L}^{k}$ represents the $k$-th street view location feature in the same region. The variable $m$ represents the count of street views and their corresponding locations within the region. $\operatorname{Aggr}(\cdot)$ denotes the aggregation method for all street views in the region, utilizing additive aggregation in this case. The function f  signifies the feature fusion method.
To establish a global-level semantic correspondence between satellite images and their related descriptions, we employ a joint optimization approach for the image and text modal encoders in the satellite branch. For the $i$-th satellite image-report pair $(I_{g_{i}},T_{g_{i}})$ in a mini-batch, we regard two modality data as queries and keys alternatively to learn the correct image-report pairings, to maximally preserve the mutual information between the true pairs in latent space. The global contrastive loss function $\mathcal{L}_{CG}$ is composed of two terms: $\mathcal{L}_{{CG}}^{Image\rightarrow Text}$ and $\mathcal{L}_{{CG}}^{Text\rightarrow Image}$, which measure the similarity between the image and text embeddings, respectively, defined as:
(4) $\mathcal{L}_{{CG}}^{Image\rightarrow Text}=-\sum_{i}^{N}{\log\frac{\exp(I_{g_{i}}^{\top}T_{g_{i}}/\tau)}{\sum_{j=0}^{N}{\exp(I_{g_{i}}^{\top}T_{g_{j}}/\tau)}}},$
(5) $\mathcal{L}_{{CG}}^{Text\rightarrow Image}=-\sum_{i}^{N}{\log\frac{\exp(T_{g_{i}}^{\top}I_{g_{i}}/\tau)}{\sum_{j=0}^{N}{\exp(T_{g_{i}}^{\top}I_{g_{j}}/\tau)}}},$
(6) $\mathcal{L}_{CG}=\frac{1}{N}(\mathcal{L}_{{CG}}^{Image\rightarrow Text}+\mathcal{L}_{{CG}}^{Text\rightarrow Image}),$
where $I_{g_{i}}$ and $T_{g_{j}}$ are the normalized embedding of satellite image representation $\mathbf{z}_{g_{sa}}$ in the $i$-th pair and that of text representation $\mathbf{z}_{T_{sa}}$ in the $j$-th pair, respectively. Besides, $N$ is the batch size, and $\tau$ is the temperature for contrastive learning.
For capturing local information, given that the street-view local branch is expected to capture fine-grained information to provide ground-level details for region embedding, approaches like (Radford et al., 2021; Jia et al., 2021) rely exclusively on the similarity of each modality’s global feature, thus lacking the capability to capture finer-level information such as the relationship between visual objects and textual tokens (Yao et al., 2022a). We utilize a fine-grained interaction mechanism to implement alignment. Specifically, we utilize token-level maximum similarity between visual and textual tokens to direct the contrastive objective. Initially, we compute the similarity between each visual token and all textual tokens, and then leverage the maximum value to calculate the average similarity of all image tokens to textual tokens. The similar approach is also applied to text-to-image.
(7) $\operatorname{SIM}(v_{i},t_{i})=\frac{1}{l_{1}}(\sum_{k_{1}=1}^{l_{1}}v_{ik_{1}}^{\top}{\operatorname{argmax}_{k_{2}\in[0,l_{2})}(t_{ik_{2}})}),$
(8) $\operatorname{SIM}(t_{i},v_{i})=\frac{1}{l_{2}}(\sum_{k_{2}=1}^{l_{1}}t_{ik_{1}}^{\top}{\operatorname{argmax}_{k_{2}\in[0,l_{1})}(v_{ik_{2}})}),$
where $v_{i}$ denotes the $i$-th token of visual features and $t_{j}$ represents the $j$-th token of textual features. $l_{1}$ and $l_{2}$ denote the token length of visual and textual features in street-view branch respectively. The fine-grained token-level representation can be optimized through:
(9) $\mathcal{L}_{{CL}}^{Image\rightarrow Text}=-\sum_{i}^{N}{\log\frac{\exp(\operatorname{SIM}(v_{i}^{\top},t_{i})/\tau)}{\sum_{j=0}^{N}{\exp(\operatorname{SIM}(v_{i}^{\top},t_{j})/\tau)}}},$
(10) $\mathcal{L}_{{CL}}^{Text\rightarrow Image}=-\sum_{i}^{N}{\log\frac{\exp(\operatorname{SIM}(t_{i}^{\top},v_{i})/\tau)}{\sum_{j=0}^{N}{\exp(\operatorname{SIM}(t_{i}^{\top},v_{j})/\tau)}}},$
(11) $\mathcal{L}_{CL}=\frac{1}{N}(\mathcal{L}_{{CL}}^{Image\rightarrow Text}+\mathcal{L}_{{CL}}^{Text\rightarrow Image}).$
3.3. Urban Indicator Prediction
Pre-training Stage. The total objective of UrbanVLP can be defined as the sum of the above two losses:
(12) $\mathcal{L}_{Total}=\alpha\mathcal{L}_{{CG}}+\beta\mathcal{L}_{{CL}}.$
where $\alpha$ and $\beta$ are weight hyperparameter. Through backpropagation optimization, We achieve multi-granularity cross-modal urban image-text input data alignment, resulting in a robust encoder.
Prediction Stage. During the fine-tuning stage, we employ a linear probing approach (He et al., 2022). This process begins by extracting frozen features $\mathbf{e}_{sa},\mathbf{e}_{sv},\mathbf{e}_{t}$ from the pre-trained encoder for satellite images, street view images, and street view positions (The latter two are optional). Subsequently, these features are fused, and a minimalist classifier (MLP) is trained on top to fine-tune the prediction of urban metrics, denoted as $\mathbf{Y}_{i}=MLP(\mathbf{e}_{sa},\mathbf{e}_{sv},\mathbf{e}_{t})$. As depicted in Figure 3, these encoders have adapted to capturing high-level abstractions from extensive visual data, enabling the MLP to manipulate static features and study the model’s inference capabilities.
4. Experiments
In this section, we conduct extensive experiments to investigate the following Research Questions (RQ):
  • RQ1: Can UrbanVLP outperform previous baseline methods and generalize well to various urban indicator tasks?
  • RQ2: How does each component (e.g., satellite branch, street-view branch, location encoding branch) contribute to UrbanVLP?
  • RQ3: How do automated text generation and calibration affect the generated text quality?
  • RQ4: How do we envision the practicality of UrbanVLP?
Given the limits of page number, RQ4 is detailed in Appendix D.
4.1. Datasets & Tasks
Our dataset comprises a dual-category structure encompassing both satellite and street-view image components, each paired with corresponding textual descriptions. For the satellite image-text pairs, we follow and refine the setting of (Yan et al., 2023) to utilize core area data of 4 cities in China.
4.1.1. UrbanSV Dataset
We construct the UrbanSV Dataset by street-view geo-tagged images collected from Baidu Map API and generated text descriptions by LMM.
  • Street-View Panoramic Image Collection. The Street-View image sampling process utilized the Baidu Maps API [^3] https://lbsyun.baidu.com/, facilitating the extraction of road network data within our predefined sampling area. Using the ArcGIS [^4] https://www.esri.com/en-us/arcgis/products/develop-with-arcgis/overview tool, we exported the road network data and established a sampling interval of 500 meters, which allowed for the comprehensive collection of street-view images at every half-kilometer point along the road network, ensuring thorough coverage that complements the satellite imagery’s scope. Following the road network sampling, panoramic street-view images were collected based on the sampling points. These images, with a resolution of 2048x664, also far exceed the 256x256 resolution of satellite imagery, ensuring clarity and ample detail for further analysis. As is shown in Table 1, a total of 47,582 street-level images were collected, whose data volume is 2.7 times that of the satellite dataset, capturing the intricacies of the urban landscape and providing a street-level view that enhances the overall urban analysis.
  • Corresponding Text Description Generation. The textual data corresponding to each satellite image is generated from ShareGPT4V (Chen et al., 2023a), which is recognized for its superior text generation capabilities compared to other contemporary open-source LMM models (Liu et al., 2023a; Dai et al., 2023; Ye et al., 2023; Gao et al., 2023; Chen et al., 2023b, a), offering comprehensive and high-quality textual information. The generated quality can be guaranteed by our perception scores. We filter out the texts with PerceptionScores less than 0.5.
Table 1. Urban imagery dataset statistics.
Dataset Coverage #Street -view Image #Satellite Image
Bottom-left
Top-right
Beijing 39.75°N, 116.03°E 40.15°N, 116.79°E 14,164 4,592
Shanghai 30.98°N, 121.10°E 31.51°N, 121.80°E 15,616 5,244
Guangzhou 22.94°N, 113.10°E 23.40°N, 113.68°E 7,635 3,402
Shenzhen 22.45°N, 113.75°E 22.84°N, 114.62°E 10,167 4,324
4.1.2. Downstream Task Dataset
In our research, the Downstream Task Dataset plays a crucial role in gauging the practical implications of our multimodal framework. We collect six representative indicators: Carbon Emission, Population, GDP, Night Light, House Price, and POI. Each indicator illuminates a significant dimension of urban vitality and sustainability. The detailed introduction of each downstream indicator is provided in the Appendix A.
Table 2. Urban indicators prediction results in four datasets. The best results are in bold and the second-best results are underlined.
Methods UrbanVLP StructualUrban (Li et al., 2022c) Vision-LSTM (Huang et al., 2023a) UrbanCLIP-SV (Yan et al., 2023) UrbanCLIP (Yan et al., 2023) PG-SimCLR (Xi et al., 2022) ResNet-18 (He et al., 2016) Autoencoder (Kramer, 1991)
Metric $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE
Beijing Carbon 0.787 0.457 0.353 0.745 0.500 0.397 0.692 0.502 0.421 0.504 0.697 0.536 0.716 0.528 0.528 0.430 0.797 0.632 0.383 0.821 0.598 0.294 0.851 0.572
Population 0.725 0.513 0.404 0.699 0.536 0.422 0.649 0.583 0.511 0.447 0.727 0.574 0.661 0.569 0.452 0.476 1.228 0.963 0.251 0.874 0.657 0.164 0.918 0.673
GDP 0.586 0.650 0.416 0.560 0.670 0.418 0.536 0.690 0.454 0.208 0.899 0.553 0.555 0.674 0.439 0.270 1.679 1.067 0.242 1.253 0.733 0.165 1.521 0.838
Night Light 0.531 0.629 0.394 0.471 0.668 0.447 0.409 0.712 0.466 0.352 0.733 0.476 0.420 0.700 0.457 0.367 0.728 0.404 0.337 0.740 0.581 0.276 0.784 0.643
House Price 0.506 0.640 0.479 0.494 0.647 0.481 0.470 0.655 0.493 0.453 0.672 0.512 0.503 0.641 0.482 0.341 0.718 0.555 0.284 0.721 0.576 0.332 0.714 0.576
POI 0.319 0.706 0.383 0.277 0.712 0.385 0.236 0.718 0.388 0.227 0.724 0.410 0.253 0.731 0.388 - - - 0.209 0.747 0.415 0.164 0.759 0.427
Shanghai Carbon 0.716 0.529 0.392 0.677 0.556 0.431 0.568 0.613 0.458 0.574 0.473 0.453 0.671 0.569 0.426 0.270 0.742 0.551 0.243 0.763 0.528 0.230 0.771 0.532
Population 0.593 0.607 0.471 0.548 0.630 0.486 0.521 0.657 0.518 0.402 0.793 0.590 0.536 0.648 0.507 0.288 1.051 0.831 0.274 0.833 0.632 0.166 0.898 0.742
GDP 0.310 0.816 0.595 0.305 0.841 0.613 0.304 1.012 0.664 0.315 1.214 0.741 0.326 0.807 0.587 0.270 1.679 1.067 0.241 1.237 0.740 0.197 1.511 0.844
Night Light 0.457 0.667 0.459 0.412 0.689 0.494 0.355 0.731 0.546 0.276 0.763 0.554 0.387 0.709 0.511 0.228 0.596 0.466 0.202 0.789 0.573 0.164 0.810 0.617
POI 0.413 0.673 0.363 0.387 0.681 0.375 0.247 0.699 0.389 0.243 0.707 0.399 0.269 0.704 0.394 - - - 0.221 0.723 0.411 0.186 0.729 0.425
Guangzhou Carbon 0.698 0.514 0.385 0.654 0.550 0.506 0.551 0.564 0.462 0.423 0.520 0.478 0.585 0.603 0.444 0.413 0.951 0.716 0.362 0.673 0.536 0.254 0.733 0.570
Population 0.665 0.565 0.441 0.638 0.586 0.453 0.617 0.615 0.506 0.497 0.685 0.612 0.633 0.592 0.467 0.294 1.046 2.112 0.249 0.834 0.678 0.195 0.928 0.753
GDP 0.436 0.764 0.541 0.419 0.791 0.549 0.405 1.011 0.625 0.321 1.004 0.689 0.440 0.762 0.546 0.263 2.162 1.553 0.234 1.147 0.732 0.176 1.463 0.811
Night Light 0.577 0.573 0.418 0.534 0.596 0.442 0.405 1.011 0.625 0.335 0.672 0.514 0.483 0.633 0.478 0.436 0.549 0.427 0.225 0.720 0.562 0.176 0.798 0.620
POI 0.396 0.752 0.344 0.325 0.771 0.364 0.143 0.822 0.411 0.136 0.823 0.415 0.150 0.819 0.408 - - - 0.108 0.920 0.433 0.084 0.956 0.469
Shenzhen Carbon 0.659 0.568 0.418 0.613 0.570 0.434 0.564 0.576 0.471 0.483 0.586 0.488 0.562 0.571 0.483 0.234 0.975 0.746 0.207 0.735 0.597 0.189 0.887 0.634
Population 0.790 0.448 0.343 0.724 0.476 0.384 0.624 0.562 0.471 0.589 0.604 0.483 0.727 0.510 0.392 0.294 1.046 2.112 0.264 0.871 0.694 0.175 0.921 0.772
GDP 0.532 0.676 0.448 0.489 0.695 0.471 0.462 0.720 0.517 0.433 0.754 0.577 0.508 0.693 0.464 0.294 1.046 2.112 0.182 1.245 0.836 0.119 1.733 0.884
Night Light 0.457 0.667 0.459 0.421 0.694 0.502 0.404 0.702 0.508 0.320 0.717 0.531 0.387 0.709 0.511 0.454 0.199 0.158 0.243 0.726 0.543 0.166 0.758 0.608
POI 0.461 0.759 0.347 0.321 0.780 0.374 0.254 0.795 0.396 0.147 0.849 0.440 0.185 0.838 0.427 - - - 0.126 0.968 0.469 0.092 0.973 0.479
$1^{\text{st}}$ount 53 0 0 1 5 3 0 0
Table 3. Ablation study on UrbanSV Beijing dataset. The best results are in bold.
Model Carbon Population GDP Night Light House Price POI
$\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE
w/o SV branch 0.754 0.491 0.387 0.690 0.545 0.428 0.547 0.680 0.433 0.468 0.670 0.445 0.520 0.635 0.465 0.263 0.714 0.387
w/o ST branch 0.518 0.687 0.512 0.465 0.715 0.559 0.229 0.887 0.543 0.423 0.698 0.423 0.295 0.764 0.575 0.265 0.716 0.390
w/o LE branch 0.776 0.469 0.358 0.717 0.520 0.414 0.589 0.652 0.409 0.518 0.633 0.400 0.508 0.639 0.472 0.301 0.710 0.385
w/o SV&ST branch 0.716 0.528 0.324 0.626 0.598 0.425 0.258 0.871 0.511 0.482 0.661 0.402 0.493 0.654 0.513 0.259 0.721 0.395
UrbanVLP 0.788 0.455 0.455 0.720 0.517 0.410 0.585 0.651 0.420 0.523 0.634 0.398 0.523 0.634 0.407 0.319 0.706 0.383
4.2. Experimental Setup
4.2.1. Baselines
We compare our method with 7 baselines in the field of urban imagery-based socioeconomic prediction. The single granularity imagery-based urban region representation learning methods include: AutoEncoder (Kramer, 1991), ResNet-18 (He et al., 2016), PG-SimCLR (Xi et al., 2022), UrbanCLIP (Yan et al., 2023), UrbanCLIP-SV (Yan et al., 2023). We also apply multiple baselines for multi-granularity urban imagery-based representation learning: Vision-LSTM (Huang et al., 2023b) and StructualUrban
(Li et al., 2022c)
. The in-depth overview of each baseline method is presented in the Appendix B.
4.2.2. Metrics and Implementation
The performance of our predictions is evaluated using a trio of standard metrics: the coefficient of determination denoted as $R^{2}$, the root mean squared error (RMSE), and the mean absolute error (MAE). An increase in $R^{2}$, along with a reduction in RMSE and MAE values, is indicative of enhanced model accuracy. Adam optimizer is chosen to minimize the training loss during parameter learning. A grid search on hyperparameters is conducted, where search ranges for learning rate and batch size are set as 2e-6, 2e-5, 2e-4, 2e-3, 2e-2 and 4, 8, 16, 32, 64, respectively. The data augmentation strategy follows the setting of (Radford et al., 2021). The weight hyperparameter $\alpha$ and $\beta$ of $\mathcal{L}_{Total}$ are set to 0.5, 0.5 respectively. We run all the models on NVIDIA A800 GPUs.
We implement the aggregation method $Aggr$ through MLP layers. To ensure a fixed length for network processing, we set the max street-view sequence length within one as 25, and sequences shorter than that are achieved by padding with zeros. The fusion method is Addition. We utilize the current state-of-the-art GeoCLIP (Vivanco et al., 2023) as our location encoder. For the semantic segmentation model to decouple visual elements with street-view images, the architecture for the semantic model is ResNet18dilated + PPM_deepsup. The original dataset contains 150 categories, from which we select 38 categories and group them into 13 street-view features following. (Fan et al., 2023b). The detailed categories can be seen in Appendix C.
4.3. RQ1: Performance Evaluation
To further validate the effectiveness of the proposed UrbanVLP, we conduct comparisons with existing state-of-the-art methods using all datasets. The comprehensive results are presented in Table 2. As we can see, UrbanVLP significantly surpasses the performance of baseline methods employing both single-granularity and multi-granularity imagery, achieving a substantial margin of improvement. It outperforms the best baseline, StructualUrban (Li et al., 2022c), by 4.2%, 2.6%, 2.6%, 6.0%, 1.2% and 4.7% in terms of $R^{2}$ for all 6 indicators: Carbon, Population, GDP, Night Light, House Price and POI in Beijing. The similar performance improvement can also be seen in the dataset of the other three cities. Besides, the average performance gain of UrbanVLP on RMSE and MAE in Beijing data are 2.3% and 2.0%, respectively. In order to provide a more intuitive display, we have created radar Image in Figure 2 to visualize the performance of two cities, Beijing and Shenzhen. The results further prove the effectiveness of our multi-granularity framework for urban region representation learning.
It is evident that pipelines leveraging multi-granularity approaches typically demonstrate superior performance compared to those employing a single granularity level. This enhancement can be rationalized by the incorporation of additional fine-grained information derived from street-view imagery. UrbanVLP exhibits superior performance over StructuralUrban (Li et al., 2022c) and Vision-LSTM (Huang et al., 2023b), illustrating the efficacy of textual mobility, which enhances interpretability. Concurrently, it is observed that UrbanCLIP-SV significantly underperforms compared to UrbanCLIP, demonstrating that street-view imagery, while rich in fine-grained details, lacks macroscopic perspective, leading to suboptimal outcomes when used in isolation. This also underscores the irreplaceability of satellite imagery as a macroscopic visual modality.
4.4. RQ2: Ablation Studies
We conduct ablation studies to investigate the effectiveness of different components in UrbanVLP on UrbanSV Beijing dataset, including the Satellite branch (ST branch), Street-View branch (SV branch), and Location Encoding branch (LE branch). The results are demonstrated in Table 3. It can be observed that the incorporation of street-view imagery results in an average improvement of 3.75% in terms of $R^{2}$. Furthermore, while street-view imagery contributes fine-grained information, its individual impact is still outstripped by that of satellite imagery, aligning with the performance differential between UrbanCLIP and UrbanCLIP-SV in Section 4.3. The role of location coordinates has been underexplored in previous work. Although not as impactful as visual modalities, their inclusion has yielded an average improvement of 0.5% in terms of $R^{2}$.
In our work, we leverage ShareGPT4V (Chen et al., 2023a) for text description generation. Here we conduct experiments on current open-sourced LMM models, and compute the PerceptionScore of them. The results are presented in Table LABEL:table:lmms. As we can see, ShareGPT4V achieves the best performance in terms of PerceptionScore, followed by LLaMA-Adapter V2, which are used as the Image-to-Text models (Yan et al., 2023).
Table 4. Perception score of representative LMMs.
Models Average Generated-Description Scores
LLaVA-v1.5 (Liu et al., 2023a) 0.674
mPLUG-Owl (Ye et al., 2023) 0.662
LLaMA-Adapter V2 (Gao et al., 2023) 0.698
MiniGPTv2 (Chen et al., 2023b) 0.693
ShareGPT4V (Chen et al., 2023a) 0.714
Figure 5. Illustration of the quality of our generated description example. The prompt for (c) is :”From scale 1 to 10, decide how similar the image is to the description”, whereas the prompt for (d) is: ”Generate a panoramic street-view image based on the description”.
Figure 6. Perception Score distribution on UrbanSV dataset for four representative cities.
4.5. RQ3: Qualitative Analysis
To visually illustrate the quality of the generated description, we present an example in Figure 5. The generated description of the street-view image (a) is depicted in (b). Subsequently, we employ GPT4V (Yang et al., 2023) to assess the quality score of the generated description, which is rated as 7 out of 10, indicating the effectiveness of the generated text. Furthermore, GPT4V also provides specific areas for improvement in (c). In (d), we utilize GPT4V to generate an image based on the description in (b). It is evident that the generated image bears a resemblance to the original one. In order to ascertain the efficacy of ShareGPT4V, we computed the distribution of PerceptionScores for descriptions generated by ShareGPT4V, as depicted in Figure 6. Notably, only less than 0.02% of image-text pairs scored below 0.5, thereby evidencing the high quality of the descriptions produced.
Figure 7. Examples of predictions from UrbanCLIP and our proposed UrbanVLP.
Here we also show some predicted results in Figure 7 on UrbanSV Beijing dataset. As we can see, satellite image (a) and (b), despite sharing similar layouts, (a) encompasses residential and campus areas, while (b) pertains to industrial zones. Their socio-economic characteristics are entirely different, which could be reflected by socio-enconomic indicators. The carbon emissions and GDP of campus and residential area are significantly lower than those of industrial park, whereas the disparities in population and housing prices are not as pronounced, possibly due to the unique characteristics of school district housing near educational institutions. Night light intensity is similar across the two area due to the fact that the majority of nocturnal economic activities within factories occur indoors, unlike in Central Business Districts (CBDs) where extensive illumination is generated. The same principle applies to school and residential area as well. As observed, the predictions made by UrbanCLIP do not effectively differentiate between the two, with the downstream metrics in Figure 7 (b) still inclined towards predicting a homogenized socio-economic functional area similar to that of (a). In contrast, the results from UrbanVLP are capable of distinctly distinguishing the socio-economic attributes of the two in terms of carbon emission, GDP, and house price.
5. Conclusion and future work
Urban region representation learning plays a significant role in helping us understand the patterns, structures, and dynamics of cities. UrbanVLP, for the first time, explores the differences between street-view and satellite images from a semantic granularity perspective, as well as their roles in modeling urban region representation. A text generation and calibration mechanism is also proposed to ensure high-quality description generation. It has achieved state-of-the-art results on the constructed UrbanSV and downstream dataset. In the future, the further research direction can be two-fold: Add additional modalities such as POIs to enhance information richness and enable the design of better model architectures such as MAE (He et al., 2022) to efficiently utilize existing data.
Appendix A Detailed introduction of downstream dataset
We provide the detailed introduction and data source of our downstream benchmark.
  • Carbon Emissions: Extracted from the Open-source Data Inventory for Anthropogenic CO2 (ODIAC), this dataset offers granular carbon emission figures in 2022 that are in alignment with the spatial resolution of our satellite imagery, which is one square kilometer per image. The unit is ton.
  • Population: Sourced from WorldPop, this data provides population distribution data in 2020. The unit is #citizens.
  • GDP: This data reflects the economic development situation in China in terms of Gross Domestic Product (GDP) figures to add an economic layer to our urban region analysis.
  • Night Light: Nighttime light can, to some extent, characterize the intensity of human activities, which is of great significance for urban development studies. we take the nightlight data from (Zhong et al., 2022) in 2020.
  • House Price: The price of housing is influenced by a variety of socio-economic factors. By observing changes in housing prices, it is possible to discern, to a certain extent, the state of economic development. We collect data from Anjuke (https://www.anjuke.com) in 2023.3. The unit is yuan/$m^{2}$. We take the average of housing prices within an urban region.
  • POI: The distribution of different types of POIs (Points of Interest) within urban areas can serve as a social indicator, as they reflect the social and economic structure, population activities, and lifestyles of the city. We collect POIs (location and category) from the web. The data has 14 categories: Dining and Cuisine, Leisure and Entertainment, Sports and Fitness, Business and Residential, Healthcare, Financial Institutions, Tourist Attractions, Lifestyle Services, Shopping and Consumption, Automobile Related, Hotel Accommodation, Transport Facilities, Science, Education and Culture, Companies and Enterprises. We take aggregation of the number of POIs of each category within each region.
We construct this benchmark to comprehensively evaluate the learned urban region representation. We uniformly partition the downstream dataset into train:val:test=6:2:2. These indicators constitute the downstream task dataset, covering economic (GDP, Night Light, House Price), social (Population, POI), and environmental (Carbon Emission) indicators. The data preprocessing setting adheres to (Yan et al., 2023) with refinement.
Appendix B In-depth overview of baseline methods
The single granularity imagery-based urban region representation learning methods include:
  • AutoEncoder (Kramer, 1991): An architectural framework for neural networks that is designed to extract representations from unlabeled satellite imagery with reconstruction loss.
  • ResNet-18 (He et al., 2016): The well-known residual neural network pre-trained on ImageNet (Deng et al., 2009), which can directly extract visual features of satellite imagery with the knowledge acquired from natural imagery.
  • PG-SimCLR (Xi et al., 2022): A contrastive learning framework which introduces geographic information (i.e., POI) into urban region representation learning. We don’t use PG-SimCLR to predict POI for a fair comparison in our study.
  • UrbanCLIP (Yan et al., 2023): A model that utilizes an image-text contrastive learning-based approach for learning robust representations from urban imagery, intending to capture the complexity and diversity of urban areas.
  • UrbanCLIP-SV (Yan et al., 2023): Original UrbanCLIP is pre-trained on satellite image-text pairs. Here we also verify its performance on our proposed UrbanSV dataset, utilizing the same aggregation method to acquire regional features.
We also apply multiple baselines for multi-granularity urban imagery-based representation learning:
  • Vision-LSTM (Huang et al., 2023b): A vision long short-term memory neural network which could process variable-length street-view image within each region. The integration of satellite and street-view imagery enables a more effective capture of comprehensive urban representation, achieving reliable recognition results.
  • StructualUrban(Li et al., 2022c): The graph-based framework for urban region profiling leveraging street segments as containers to adaptively fuse the features of multi-level urban images. We term this method StructualUrban for simplicity.
Appendix C Segmentation Ratio and Prompt Examples
As illustrated in Figure 8, we demonstrate examples of the segmentation ratio of street-view images in our UrbanSV dataset. The segmentation ratio is decoupled by pre-trained segmentation models and is used in both prompt construction and perception score computation. We group the segmentation categories into 13 categories following (Fan et al., 2023b): Light Vehicle; Façade; Window & Opening; Road; Sidewalk; Street Furniture; Greenery - Tree; Greenery - Grass & Shrubs; Sky; Nature. Figure 9 illustrates some exmaples of our prompts used for description generation.
Figure 8. Segmentation Ratio.
Figure 9. Prompt Examples.
Appendix D RQ4: Practicality
Figure 10. UrbanVLP system.
To demonstrate the practical applications of our model in real-life scenarios, we have developed a system bearing the same name UrbanVLP. The UrbanVLP system is a novel web-based system developed to integrate LMMs with a multi-scale urban indicator framework using the Mapbox platform (Mapbox, [n. d.]). This system renders a visual representation of urban areas through satellite imagery, complemented by street-level photographs for an interactive exploration experience. Figure 10 showcases the user-friendly interface of the map, highlighting functionalities that facilitate zooming, searching for locations, and navigating through various sectors.
Each urban region captured by a satellite image, encompassing a 1km by 1km area, is marked by a blue dot, which is encircled by yellow points that denote the locations of street-view images accessible within that region. Engaging with these points allows users to uncover a suite of urban metrics, such as carbon emissions, population, GDP, night light, house price, and POI. The system’s visual components are augmented by a captioning feature that provides clear and informative textual descriptions, offering insights into the spatial characteristics of the designated area.
Furthermore, a street-view panel allows users to peruse individual street-view images alongside pertinent captions. The system also facilitates the discovery of popular POIs within any given area, enriching the understanding of its functional characteristics. Ultimately, UrbanVLP is designed to distill complex urban data into a visual and user-friendly format, thereby offering a holistic and detailed view of urban landscapes and their key indicators.